OpenStudy (amtran_bus):

Solve this trig identity (tan-1) (sec-1) =0

OpenStudy (amtran_bus):

I know sec is 1/cos But I don't think it will help.

OpenStudy (amtran_bus):

Could I square both sides and use the identity that sec^2 -1 =tan?

OpenStudy (saifoo.khan):

\[(\frac{\sin x}{\cos x} -1) ( \frac{1}{\cos x}-1)\]

OpenStudy (amtran_bus):

Then cross multiply?

OpenStudy (saifoo.khan):

You can use the FOIL method, or take LCM and solve each parenthesis first.

OpenStudy (amtran_bus):

Ok, can you show me how to foil it. I attempted it.

OpenStudy (saifoo.khan):

\[\frac{\sin x }{\cos x \times \cos x} - \frac{sinx x}{\cos x} - \frac{1}{\cos x} + 1\]

OpenStudy (saifoo.khan):

This will lead you to: \[\Large \frac{\sin x}{\cos^2 x} - \frac{\sin x-1}{\cos x} + 1\]